OpenStudy (anonymous):

Find the derivative of the function. MEDAL WILL BE REWARDED! g(x)=sqrt(5-3x)

hartnn (hartnn):

chain rule know what that is ?

OpenStudy (anonymous):

Yes what should I do? Im not 100% sure on how to use chain rule @hartnn

hartnn (hartnn):

what is the derivative of sqrt x ?

OpenStudy (anonymous):

1/2x^2?

hartnn (hartnn):

1/2 sqrt x isn't it ?

OpenStudy (anonymous):

wait......yes that's it okay

hartnn (hartnn):

so, d/dx (5-3x) = 1/[2 sqrt (5-3x) ] times d/dx (5-3x) so whats d/dx (5-3x) ?

OpenStudy (anonymous):

5

hartnn (hartnn):

no.... derivative of constant = 0

OpenStudy (anonymous):

so wait, I don't see how that's a constant...oh wiat nvm. Okay so our derivative is 0

hartnn (hartnn):

no, d/dx (-3x) = -3

OpenStudy (anonymous):

ohh im so sorry I thought it was 5x-3 not 5-3x, okay I got -3

hartnn (hartnn):

so, d/dx sqrt (5-3x) = -3/[2 sqrt (5-3x) ]

OpenStudy (anonymous):

okay, is that my final answer? or should I simplify?

hartnn (hartnn):

no, final answer..

OpenStudy (anonymous):

okay, thank you so much!

hartnn (hartnn):

welcome ^_^